OpenStudy (krgatorgirl1):

3/5 piece of tape he wants to divide into 1/10 sections how many sections will the piece of tape be divided into

OpenStudy (leyb):

\[\frac{ 3 }{ 5 }\div \frac{ 1 }{ 10 }\] \[= \frac{ 3 }{ 5 }\times \frac{ 10 }{1 }\] = 6 There are 6 sections.

OpenStudy (leyb):

You can post your questions like this in the Mathematics section so you can get more help and explanation.

OpenStudy (krgatorgirl1):

Thanks

OpenStudy (leyb):

@Krgatorgirl1 Thank you for the medal!